OpenStudy (ggray):

What is the product in scientific notation? (2.5*10^-10)(7*10^-6)

OpenStudy (ggray):

A= 1.75*10^16 B= 1.75*10^-15 C= 1.75*10^-16 D= 1.75*10^-17

OpenStudy (ggray):

@reneezerichter I think it is b am I right?

OpenStudy (ggray):

Please?

OpenStudy (ggray):

@sissyedgar do u know if I am right?